Consider the functions

f(x)=5 and g(x)=5x.

f(x)=5 and g(x)=5x.

The value of g(x)−f(x)

g(x)−f(x) is


1
Expert's answer
2022-03-01T08:37:17-0500

Solution:

f(x)=5 and g(x)=5x

g(x)−f(x) = 5x-5 = 5(x-1)